OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. A(5, 5), m = 3

OpenStudy (texaschic101):

y = mx + b slope(m) = 3 (5,5)..x = 5 and y = 5 now we will sub and find the y intercept, which is in the b position 5 = 3(5) + b 5 = 15 + b 5 - 15 = b -10 = b so our equation is : y = 3x - 10...but we need it in standard form : Ax + By = C y = 3x - 10 --- subtract 3x from both sides -3x + y = -10 -- multiply by -1 to make x positive 3x - y = 10 <== standard form

OpenStudy (anonymous):

Thank you! @texaschic101

OpenStudy (texaschic101):

sure thing :)